### v1

# Tool-LMM: A Large Multi-Modal Model for Tool Agent Learning

Abstract
Recently, the astonishing performance of large language models (LLMs) in natural language comprehension and generation tasks triggered lots of exploration of using them as central controllers to build agent systems. Multiple studies focus on bridging the LLMs to external tools to extend the application scenarios. However, the current LLMs’ perceiving tool-use ability is limited to a single text query, which may result in ambiguity in understanding the users’ real intentions. LLMs are expected to eliminate that by perceiving the visual- or auditory-grounded instructions’ information. Therefore, in this paper, we propose Tool-LMM, a system incorporating open-source LLMs and multi-modal encoders so that the learnt LLMs can be conscious of multi-modal input instruction and then select the function-matched tool correctly. To facilitate the evaluation of the model’s capability, we collect a dataset featured by consisting of multi-modal input tools from HuggingFace. Another important feature of our dataset is that our dataset also contains multiple potential choices for the same instruction due to the existence of identical functions and synonymous functions, which provides more potential solutions for the same query. The experiments reveal that our LMM is capable of recommending appropriate tools for multi-modal instructions. Codes and data are available at https://github.com/Tool-LMM/Tool-LMM.
1 Introduction
Large Language Models (LLMs)  [2, 20, 25, 33] have showcased remarkable success in conversational scenarios. With the promising capability of the instruction following via reinforcement learning from human feedback (RLHF)  [1, 4, 13], developers extend LLMs, e.g., ChatGPT with a Function Call feature, to the tool-use domain. This extension of LLMs’ capabilities can be applied to a broad spectrum of situations, from intriguing outer tasks like restaurant reservations and online shopping to inner operations involving the retrieval and processing of infinite external knowledge. Along this direction, numerous intelligent agents  [23, 31, 24, 41, 3, 43, 7, 18] have been developed, by leveraging LLMs to accomplish various tasks in human-like ways.
Figure 1: Comparisons of results outputted by tool agent with different inputs. The individual posed the same question thrice. Initially, no image was provided, causing the confusion. In the subsequent instances, images from varied scenes were uploaded, and our Tool-LMM can correctly choose the correct tool based on the multi-modal inputs and generate distinct responses based on the differing pictorial contexts.
Although some efforts  [15, 39, 23, 21] have been made to stimulate the capability to use external tools with LLMs, they are confined to purely textual inputs, leading to potential confusion and misinterpretation, especially when text is ambiguous in specifying the users’ real intentions. As illustrated in Figure 1, for the given textual prompt, “Generate a picture conditioned on the image. A professional athlete dressed in a football uniform, throwing his arms up in a cheer.”, LLMs struggle due to a lack of visual input and cannot specify the type of conditioned image. Recognizing this as an image generation task, they falter in recommending suitable pose- or scribble-conditioned text-to-image tools. However, when endowed with the capability to perceive visual or auditory instructions, LLMs, in this instance, can extract key “human pose” and “scribble” information from the second time and third time input images, respectively, thus resolving the ambiguity issue. Subsequently, multimodal instruction sensing is necessary to identify the appropriate tools for different tasks.
Method Trainable LLM Multimodal LLM Num. of modality Multi-options
MetaTool [12] 1
APIBank [15] 1
ToolBench [39] 1
ToolLLM [23] 1
APIBench [21] 1
Visual ChatGPT [37] 2
GPT4Tool [42] 2
HuggingGPT [27] 4
Ours 4
Table 1: Comparison of previous works and our work. The first five are intended for tool-augmented LLMs with textual inputs, and the middle three focus on empowering the task planning of LLMs and executing multi-modal inputs with corresponding tasks’ APIs.
To remedy the previous works only accepting single text instructions, in this paper, we introduce a novel system, Tool-LMM, integrating multi-modal encoders with open-source LLMs to synthesize multi-modal information for correct external tool identification. The distinction between our work and previous works is delineated in Table 1. Compared to previous open-sourced works, our system is the first attempt to train a large multi-modal model for tool agent learning. In light of the fact that existing tool learning benchmarks  [15, 12, 39, 23, 21, 37, 42, 27], do not explicitly construct text-to-any-modality pairs for alleviating text ambiguity, which cannot facilitate the training of our Tool-LMM, in this paper, we collect a novel dataset, named ToolMMBench from the HuggingFace platform, one of the largest machine learning communities containing thousands of APIs. It is also worth noting that our ToolMMBench supports multi-options for each query, as demonstrated in Table 1. This feature is more closely related to real-life scenarios, where multiple alternatives exist to address the same challenges, albeit with varying performance levels. Such a feature would also make our dataset suitable for other applications in future, such as the study of integrating multiple outputs from different tools to better feedback to the user.
In summary, our contributions are as follows:
  • We develop Tool-LMM, a multi-modal tool agent system that can perceive the visual- or audio-grounded instructions’ information via incorporating multi-modal encoders and open-source large language models.
  • We construct a multi-modal input benchmark for evaluating LLMs’ awareness and selection ability of the external tool usage with text-ambiguous queries. The benchmark offers content-aware multi-modal instructions and one-to-many instruction-answer pairs.
  • We design some evaluation metrics and conduct experiments with extensive subset ablations on multiple popular LLMs, and by fine-tuning Tool-LMM on ToolMMBench, we observe a promising accuracy, 88.19% of the tool selection. which demonstrates the effectiveness of our method.
2 Related Works
Multi-modal Large Language Models. LLMs have increasingly become the cornerstone of many natural language processing (NLP) applications. Subsequently, many efforts, including GPT-4V  [19], LLaVA  [16], BLIP-2  [14], Minigpt-4  [46] developed the LLM to vision space by aligning off-the-shelf frozen pre-trained image encoders. And similar ideas are generalized to video and audio modalities, like NeXTGPT  [38], and VideoLLaMA  [44]. Our proposed system, Tool-LMM, focuses on exploring tool usage awareness and selection. Our goal is to empower the LLM’s ability to be conscious of the external tools’ function and select the suitable tools, to act as a controller, instead of enhancing the ability of multi-modal understanding and reasoning.
Tool Agent. Recent advancements have witnessed a rapid emergence of systems that deploy LLMs as a central orchestrator for tool assistance. Toolformer  [26] pioneers the study of bridging the LLMs with the external tools. Subsequent works have expanded the application domains of this integration to encompass health support  [17], code synthesis  [15, 30], web searching  [28]. Gorilla  [21] collected a machine learning domain dataset and introduced API retrieval to boost the performance. ToolLLM  [23] focused on executing complex tasks in practical scenarios. GPT4Tools  [42] and Visual ChatGPT  [37] integrating visual found models after decomposing the task. On the contrary, we concentrate on completely relying on the ability of large language models to naturally process multi-modal inputs and select the appropriate APIs. And to the best of our knowledge, we are the first to explore this direction.
3 Dataset Construction
(a) The distribution of sub-testing sets divided by API option number and ambiguity types respectively.
(b) The distribution of each ambiguity type in the whole dataset.
(c) The distribution of each ambiguity type in the testing set.
Figure 2: Dataset statistic visualization.
In this section, we present the process of constructing a benchmark from HuggingFace, which primarily involves three steps: collecting API datasets (Sec. 3.1), generating instruction-answer pairs (Sec. 3.2), and constructing one-to-many instruction matching (Sec. 3.3). We also introduce the statistics and distribution of the benchmark in Sec. 3.4.
3.1 API Collection
Initially, we crawled 263,945 models from the HuggingFace platform. However, the quality of these models varies significantly, for example, some models are poorly maintained, even lacking information in the model card. To ensure only those models meeting specific criteria (e.g. having a detailed model card, especially providing model description and example script) are ultimately selected, we employ a meticulous filtering process and the detailed filtering process is shown in the Appendix. Finally, we only retain 932 high-quality models.
For each model, we document the model card and extract the key information, including the domain, api_\_name, api_\_call, parameters, coarse_\_functionality, fine_\_functionality, example codes, and descriptions. Further, to prevent an overabundance of APIs with identical functionalities leading to excessive options for subsequent instruction matching (e.g., in the imaging realm, where two APIs might differ only in their network architectures but are identical in terms of the datasets they were trained on and their functionalities), we adopt a strategy based on the API’s download statistics. For any single functionality category, we retain a maximum of five of the most downloaded APIs.
It is worth noticing that the function classification rules formulated by the HuggingFace platform are coarse-grained and not suitable for direct use by our work for some reasons. In particular, some models lack assigned categories or are incorrectly classified, or they contain multiple tasks but only one of them is labeled. Additionally, the platform employs overly coarse, resulting in a loss of fine-grained distinctions. For example, the task of Audio Classification, can be subdivided into Event Recognition, Command Recognition, Spoken Language Identification, Speaker Verification, etc. We then propose a new hierarchical function categorization system and rearrange the filtered model. We list all the tasks our system supports in Table 2, and for the subtasks we divide under these tasks, we provide more detailed information in the Appendix.
Audio Classification Audio-to-Text Audio-to-Audio
Voice Activity Detection Video Classification Video-to-Text
Depth Estimation Image Classification Image Segmentation
Image-to-Image Visual Question Answering Image-to-Text
Zero-Shot Image Classification Multimodal-to-Image Object Detection
Feature Extraction Fill-Mask Text Generation
Text Generation Text Classification Question Answering
Sentence Similarity Summarization Text-to-Image
Text-to-Speech Text-to-Video Text-to-Text
Translation Zero-Shot Classification
Table 2: The 29 coarse-grained tasks our system supports.
Given our API collection may not cover all tasks, users in real-world scenarios might encounter situations without a suitable API. Thus, we also introduce the “Unknown” option to address this issue.
3.2 Instruction-Answer Pairs Generation
To curate the instruction-answer pairs, we leverage GPT-4 to understand the functionalities of these API calls and then generate 20 queries for each API call. Detailed prompts are provided in the Appendix.
After meticulously filtering and refining through manual review, 10 well-crafted queries will be finally selected to be associated with each API call. A significant highlight of our approach is its capability to disambiguate text prompts by learning the characteristics of other modalities when handling multi-modal inputs. We elaborate on five text ambiguity cases that need to be checked:
Data models used are from different domains: It is imperative to acknowledge the inherent disparities across data domains, either auditory or visual. Owing to these pronounced differences between domains, a unified treatment could result in significant biases in outcomes. Models need to judiciously select APIs contingent on the particular data domain at hand, instead of using a unified treatment, ignoring the bias across different domains. For instance, in image segmentation tasks, it is vital to select distinct APIs tailored for specific scenarios like indoor decorations, portraits’ fashion apparel, or medical related according to the given image.
Distribution and granularity of categories in different datasets: Within the same domain, we need to consider the covered categories and category granularity variations. Especially when doing image classification tasks, specific strategies vary for common, specialized, or rare categories.
Image Quality: The most evident instance revolves around image quality, which can be compromised by multiple factors including fog, rain, low resolution, and motion blur. When confronted with vague commands such as “I want to enhance the image quality”, analyzing the reason for the low imaging quality of the input image, and then selecting the appropriate API call for quality enhancement becomes necessary.
Input Conditions of different models: In some cases, the format of the input image can also assist in resolving text ambiguity issues. For example, ControlNet  [45] posits multiple conditions for image generation. Faced with an instruction like “Based on the provided image, generate a photo of a chef in the kitchen,” and given inputs such as a pose image or a normal map, the model must leverage the unique features distinguishing between the two types of images. This discernment facilitates the selection of the appropriate API to achieve the desired output.
Figure 3: The word cloud calculated based on the frequency of task name occurrence.
Others: There are some individual cases unique to individual modality, for instance, linguistic variations in the audio realm. Within the speech recognition task, when presented with an instruction like “In my language class, we listened to a story. I managed to record it; can you tell me its content?”, the absence of explicitly provided language information necessitates the model to first ascertain the spoken language. Only then can it aptly select the corresponding API for accurate content extraction and interpretation.
For convenience, the following defaults to five ambiguity situations, referred to as Domains, Categories, Quality, Conditions and Others.
Subsequently, nuanced manual intervention is warranted to search for prompt-matching inputs, especially in scenarios involving multi-modalities, sourced either from relevant datasets or online searches. Take the image classification task for example, one might say, “I have a photo of marine life taken during my scuba diving trip. Can you identify its category?”, it becomes crucial to provide a marine creature image input to match the context of the prompt “marine life” and “scuba diving trip”, and the corresponding image path would be saved.
We collect a total of multi-modal instruction-answer pairs, including 3798 image inputs, 103 video inputs, and 726 audio inputs, totaling 72 videos, 3777 images, and 590 audio files.
3.3 Instruction Matching
In contrast to prior works, our dataset stands out with a unique feature: it encompasses numerous API calls that perform identical or similar functions. This results in a scenario where a single instruction could be associated with several API calls within our dataset, each potentially suitable for addressing it. In step two, we approach from a singular API function perspective to construct one-to-one matched instruction-answer pairs, now necessitating integration of results from the classification system established in the first step. For APIs with identical functionalities, we can match their queries immediately. However, for groups of APIs with similar functionalities, it’s essential to determine whether a query aligns with a strict or broad functional description of each API. For instance, the request, “Can you help me generate an image of a ’golden-haired magical princess’ in the Disney style?” In this scenario, the emphasis on the “Disney style” narrows down the choice of APIs to those pre-trained on Disney-themed images. Conversely, a more general instruction like, “I’m organizing a role-playing event and need a reference image. Can you generate an image of a ’masked mysterious hero’ for me?” leaves room for multiple APIs to be potential solutions, given its lack of specificity. This process yielded a dataset where each instruction corresponds to one or, in some cases, multiple relevant API calls.
Figure 4: Our network architecture. The network could processes four distinct types of input, each uniquely color-coded. These inputs, irrespective of their modality, are first encoded using modality-specific frozen encoders. Subsequently, they are passed through a trainable projection layer, aligning them within a unified language feature space. The fine-tuned Large Language Models (LLMs), augmented with Low-Rank Adaptation (LoRA), then process a combination of the prompt and the projected embeddings, enabling the accurate predicting of corresponding API name.
3.4 Dataset Statistics
In alignment with HuggingFace’s task-specific API distributions, ToolMMBench exhibits a similar long-tailed distribution, with text-related APIs being the most abundant and video-related ones the least. ToolMMBench consists of 27 coarse-grained task categories inheriting from HuggingFace, from which we expands to 29 tasks to form the first level of our classification system, and then extending to our predefined fine-grained tasks. The overall distribution graph of our support tasks can be observed in Figure 3, where the larger the font size, the higher the frequency of occurrence. We also tally the instances of each ambiguity type in Figure 2(b), amassing a total of 2533 ambiguity pairs. Predominantly, these are concentrated within conditional cases involving image inputs, comprising 1530 pairs, since our dataset covers 11 distinct condition types of ControlNet  [45], and their various combinations, achieving a considerable quantity.
We divide our dataset into training and testing sets in an 8:2 ratio. Since our task currently focuses on a closed set of APIs, our test set cannot include any APIs that haven’t been encountered in the training set. Therefore, we approach this from the perspective of queries generated by each API. Given the presence of one-to-many query scenarios in our dataset, we implement a strategy in the training set to address these cases. Specifically, we split each one-to-many query into W distinct inputs, where W represents the number of APIs that support the given query. This approach ensures that each potential API response is individually represented, allowing the model to learn the diverse range of suitable responses for a single query. In contrast to the training sampling strategy, for the one-to-many query scenarios in our dataset, each query appears only once in the test set. Finally, the training set comprises 44,533 instructions, containing 9,245 different instructions, and the testing set totaled 2,397 instructions. The detailed distribution is illustrated in Table 3.
Settings Text Audio Image Video
Training Set(w split) 31,280 2,944 10,085 224
Training Set(w/o split) 5,563 562 3,938 82
Testing Set 1,452 164 760 21
Table 3: Distribution of training set and testing set. Note that, in the training set, one-to-many scenarios are split into multiple one-to-one instructions, so the first line represents for the number of queries after splitting while the second line calculates the distinct query number of four single modality. The last line is the distribution of the whole testing set.
Particularly for the testing set, we depict in Figure 2(a) and Figure 2(c) the distribution of API quantities across each testing subset. These subsets are constructed based on three distinct partitioning criteria: ambiguity types, API option number, and modality.
4 Tool-LMM
We introduce Tool-LMM, a method that leverages the LLM performing the role of multi-modality tool agent. The architecture of this network is depicted in Figure 4. Considering the diversity of our data, which includes images, videos, and audios, we incorporate ImageBind  [9] as our primary multimodal encoder. ImageBind  [9] is designed to create a unified embedding space that integrates six distinct modalities, utilizing the advancements in large-scale vision-language models.
For the large language model component of Tool-LMM, we select a range of leading LLMs, including Vicuna  [5], Llama  [33], Llama2  [34], and Llama2-Chat  [34]. These models are chosen for their exemplary performance in various areas, such as inference capabilities, knowledge retention, and comprehension skills. Consequently, our system is equipped with the proficiency to interpret multi-modal input instructions effectively.
In our dataset. there are four types of modality-combination input: Text, Text+Image, Text+Video, and Text+Audio. The output from the LLM includes textual responses that provide tool recommendations along with their respective model card information. For the multimodal inputs in the Image, Video, or Audio modalities, we first utilize ImageBind  [9] encoders with fixed weights for feature extraction. These features are then aligned with the LLM’s feature space through a linear projection layer. To optimize our model while minimizing the number of learnable parameters, we implement Low-Rank Adaptation (LoRA)  [11] to fine-tune an open-source LLM.
Currently, the system’s output only includes the API Name in the format {api_name: [Recommended API]}. Our future endeavors aim to expand this output to include API parameters, code execution results, and elements of Chain-of-Thought (CoT) Prompting  [35]. Additionally, the current version of our system is designed for single-turn dialogues, with the intention to extend its capabilities to multi-turn conversations in subsequent developments.
5 Evaluation Metrics
Since no prior work addresses user prompt ambiguity from the perspective of constructing text-to-any-modality pairs, there are no established metrics for measuring performance. In this paper, we propose evaluation metrics tailored to the dataset’s unique attributes, ambiguity types, multiple options, and diverse modality inputs to effectively measure performance disparities.
  • Ambiguity types: In light of the five distinct ambiguity scenarios delineated above, we construct corresponding testing subsets. These facilitate the evaluation of the model’s capability to leverage multimodal information under varying types of ambiguity, providing a nuanced understanding of its performance.
  • Multiple options: Since our model has the capacity to support multiple options, we divide the instruction set into two subsets, solved only by one single API versus those allowing for multiple API options, and we expect to examine the performance gaps within two subsets.
  • Individual modality: For each modality, we partition subsets to evaluate model performance with fixed modality combinations, ensuring a comprehensive assessment across varying input configurations.
The testing set is divided into several subsets based on the three attributes. For a subset $\mathbf{S}$, supposing containing $n$ instructions $\{x_{1},x_{2},x_{3}\dots x_{n}\}$, where each instruction $x_{i}$ matches an API list $y_{i}$, which consists of a list of $M$ suitable APIs $A_{i}^{1},A_{i}^{2},A_{i}^{3},\dots,A_{i}^{M}]$. Thus, the collection of API lists corresponding to this instruction subset combined to a set $\mathbf{Y}$, $\{y_{1},y_{2},y_{3}\dots y_{n}\}$. With these symbols defined, assuming the subset $\mathbf{S}$ yields an API set $\{z_{1},z_{2},z_{3}\dots z_{n}\}$, output from our model, we can represent the accuracy of each mentioned subset using the following formula:
$\text{Acc}=\sum\limits_{i=1}^{n}\mathbbm{1}_{z_{i}\in y_{i}}/n$ (1)
The above formula also applies to the accuracy calculation for the entire testing set. Besides that, the hallucination rate is used as a metric to evaluate the performance of each testing subset and the overall testing set. For the calculation of the hallucination rate, we first assume the corpus C, containing all APIs. Hallucination refers to the phenomenon where the model outputs APIs are not present in the corpus C, essentially fabrications of the model. Taking the above subset $\mathbf{S}$ as an example, the calculation of the hallucination rate can be expressed with a formula:
$\text{Hallu}=\sum\limits_{i=1}^{n}\mathbbm{1}_{z_{i}\notin C}/n$ (2)
Besides that, unlike simple one-to-one mapping relationships, using single-instance accuracy is insufficient for evaluating the system’s comprehension of the relationships of APIs, especially in one-to-many mapping relationships. Hence, we propose utilizing recall, allowing the system to perform repeated inferences. By examining the proportion of ground truth encompassed by the collective results of multiple inference results, marked as $R_{i}^{j}$, where $i$ is the $i^{th}$ instruction in the testing set and $j$ represents for the $j^{th}$ inference. The corresponding metric calculation is expressed as follows:
$\small\text{Recall}=\frac{1}{n}\sum_{i=1}^{n}\frac{(f_{c}(Set({R_{i}^{1},R_{i}^{2}\dots R_{i}^{K}|R_{i}^{j}\in y_{i},j\in[1,K]}))}{f_{c}(y_{i})}$
where $K$ is the number of inference, $f_{c}$ is the counting function.
Additionally, to assess whether the system’s output is intuitive and whether the output format meets user needs, we introduce a metric named format accuracy. Specifically, we aim for the output to be presented in JSON format, with clearly delineated key-value pairs. In this paper, the format should be {api_name: [Recommended API]}. Any alterations to this format, including missing symbols or changing the order of placement, are all considered as formatting errors. This structure is designed to facilitate user navigation, allowing for easy selection of desired sections.
6 Experiments
In this section, we conduct experiments on our collected dataset, evaluating the performance of our system Tool-LMM. We start by introducing the experiment setup in Sec. 6.1 and data processing in Sec. 6.2. In Sec. 6.3, we assess Tool-LMM’s ability on our benchmark, followed by results analysis in Sec. 6.4. And we conduct qualitative experiment and visualize the result in Sec. 6.5.
6.1 Experiment Setup
We list the setup, including some hyperparameters used to train our model in Table 4. The setup follows  [29] and adjusts based on our computation resources(4 $\times$ A40 40G GPUs). Especially, given the larger parameter size of the 13B model, to enhance experimental performance, we increase the epoch to 10.
Configuration 7B-model 13B-model
Optimizer AdamW AdamW
Optimizer Momentum $\beta_{1}=0.9,\beta_{2}=0.95$ $\beta_{1}=0.9,\beta_{2}=0.95$
Peak learning rate 5e-4 5e-4
Weight decay 0.001 0.001
Warmup steps 10 10
Batch size 64 32
Micro-batch size 4 1
Gradient accumulation steps 4 8
Maximum target length 512 512
ImageBind Checkpoint ImageBind-Huge ImageBind-Huge
Epochs 5 10
LoRA attention dimension (r) 32 32
LoRA scaling alpha ($\alpha$) 32 32
LoRA drop out 0.1 0.1
Table 4: 7B and 13B model training configuration.
6.2 Data processing
For multimodal inputs, we adopt ImageBind’s  [9] data processing approach. For video, we evenly split into five clips and randomly sample 2 frames from each clip to cover the full length, similar to other Vision Transformer(ViT)  [6] -based works  [32, 8] with video input. For audio, we sample each input audio at 16KHz. Subsequently, we capture a log mel spectrogram featuring 128 frequency bins, following  [10]. Then, the spectrogram can be seen as a 2D image. In this way, we can use ViT for these three multimodal encoders. A modality-specific linear projection head is added to each modality’s encoder to align the feature into a uniform 1024-sized dimension.
[…]
Vicuna-7B (3epochs) 77.35 4.13 100.00
Llama-7B (3epochs) 76.64 2.80 100.00
Llama2-7B (3epochs) 76.39 5.13 100.00
Llama2-Chat-7B (3epochs) 76.18 4.92 100.00
Vicuna-7B (5epochs) 83.15 1.63 100.00
Llama-7B (5epochs) 84.48 0.58 100.00
Llama2-7B (5epochs) 80.68 1.96 100.00
Llama2-Chat-7B (5epochs) 83.60 1.34 100.00
Vicuna-13B (3epochs) 74.34 3.63 100.00
Llama-13B (3epochs) 75.18 3.21 100.00
Llama2-13B (3epochs) 64.71 3.38 100.00
Llama2-Chat-13B (3epochs) 77.51 3.38 100.00
Vicuna-13B (5epochs) 80.85 1.63 100.00
Llama-13B (5epochs) 81.44 2.00 100.00
Llama2-13B (5epochs) 67.58 1.34 100.00
Llama2-Chat-13B (5epochs) 81.35 1.42 100.00
Vicuna-13B (10epochs) 88.19 0.17 100.00
Llama-13B (10epochs) 87.86 0.17 100.00
Llama2-13B (10epochs) 81.81 0.17 100.00
Llama2-Chat-13B (10epochs) 87.78 0.08 100.00
Table 5: Experimental results of ToolLMM. We present the performance of eight large language models with different configurations, and we report the performance at different epochs of training to reflect the changing trend.
Model Domains Categories Quality Conditions Others With ambiguity Without ambiguity
[…]
Vicuna-7B (3epochs) 60.92 0.00 64.71 3.92 97.06 0.00 87.86 0.00 97.92 0.00 82.74 0.38 74.89 3.49
Llama-7B (3epochs) 54.02 1.15 66.67 0.00 47.06 0.00 87.54 0.00 100.00 0.00 78.61 0.19 76.98 5.26
Llama2-7B (3epochs) 58.62 0.00 64.71 0.00 61.76 0.00 87.86 0.32 97.92 0.00 80.11 0.19 73.50 4.02
Llama2-Chat-7B (3epochs) 66.67 0.00 74.51 0.00 79.41 0.00 88.50 0.00 100.00 0.00 84.05 0.00 73.93 6.33
Vicuna-7B (5epochs) 83.91 2.30 72.55 0.00 100.00 0.00 94.25 0.00 100.00 0.00 91.37 0.94 80.79 1.98
Llama-7B (5epochs) 75.86 0.00 66.67 0.00 73.53 0.00 95.21 0.00 100.00 0.00 88.37 0.00 83.37 0.75
Llama2-7B (5epochs) 66.67 0.00 66.67 0.00 94.12 0.00 92.97 0.00 100.00 0.00 86.87 0.00 78.92 2.52
Llama2-Chat-7B (5epochs) 87.36 0.00 74.51 0.00 97.06 0.00 95.85 0.00 100.00 0.00 92.87 0.00 80.95 1.72
Vicuna-13B (3epochs) 56.32 0.00 64.71 0.00 82.35 0.00 88.82 0.00 100.00 0.00 81.80 0.00 73.28 4.13
Llama-13B (3epochs) 24.14 0.00 66.67 0.00 44.12 11.76 89.46 0.00 100.00 0.00 74.67 0.75 74.25 4.45
Llama2-13B (3epochs) 59.77 0.00 76.47 0.00 58.82 0.00 90.10 0.00 97.92 0.00 82.55 0.00 76.07 4.35
Llama2-Chat-13B (3epochs) 39.08 0.00 68.63 0.00 38.24 0.00 26.20 1.60 97.92 0.00 39.59 0.94 71.89 4.08
Vicuna-13B (5epochs) 82.76 0.00 68.63 7.84 97.06 0.00 93.61 0.32 100.00 0.00 90.24 0.94 78.92 2.31
Llama-13B (5epochs) 48.28 0.00 62.75 0.00 100.00 0.00 88.50 0.96 100.00 0.00 81.24 0.56 80.74 1.93
Llama2-13B (5epochs) 74.71 0.00 60.78 0.00 91.18 0.00 88.82 0.00 100.00 0.00 84.99 0.00 80.31 1.82
Llama2-Chat-13B (5epochs) 33.33 0.00 58.82 0.00 41.18 0.00 31.63 0.00 87.50 0.00 40.15 0.00 75.43 1.72
Vicuna-13B (10epochs) 88.51 0.00 66.67 0.00 97.06 0.00 97.76 0.00 100.00 0.00 93.43 0.00 86.70 0.21
Llama-13B (10epochs) 89.66 0.00 76.47 0.00 100.00 0.00 97.12 0.00 100.00 0.00 94.37 0.00 86.00 0.21
Llama2-13B (10epochs) 64.37 0.00 60.78 0.00 73.53 0.00 73.48 0.32 100.00 0.00 73.17 0.19 84.28 0.16
Llama2-Chat-13B (10epochs) 90.80 0.00 78.43 0.00 97.06 0.00 97.44 0.00 100.00 0.00 94.75 0.00 85.78 0.11
Table 6: The performance on testing subsets divided by ambiguity type.
6.3 Experimental Results
We finetune four open-source LLMs, Llama, Llama2, Llama2-Chat, and Vicuna, each with 7B and 13B two variants, on the instruction-solution data pairs. As illustrated in Table 5, we report the format accuracy, the accuracy, and the hallucination rate, which are defined in Table 5 in detail. Due to the limitation of computing resources, the experiment results for recall metric, we focused on the two models, Llama2-7B and Llama2-Chat-7B, and set the inference times $K$ 10. The results are shown in Table 7.
[…]
Llama2-7B(3 epochs) 55.85
Llama2-7B(5 epochs) 62.13
Llama2-Chat-7B(3 epochs) 55.23
Llama2-Chat-7B(5 epochs) 61.47
Table 7: The recall performance of two 7B models.
We can derive some findings from the results:
1. For the 7B models, Llama-7B performs the best performance both on accuracy and hallucination rate, with 84.48% and 0.58% respectively when trained with 5 epochs. As we increase the training epochs, we can observe that there is a consistent improvement in both accuracy(+ 6.34%) and the reduction of hallucination rates(- 2.87%) averagely compared to that with 3 epochs. As for 13B models, when training for 5 epochs, the performance is not as good as the corresponding 7B model. But if increasing to 10 epochs, we can observe a substantial improvement, and Vicuna-13B achieves 88.19% accuracy. Meanwhile, all the 13B models’ hallucination rate drop to within 0.17% at 10 epochs.
2. The format maintains a hundred percent accuracy across all models, demonstrating LLMs have a high understanding of structural rules, which is crucial for tasks that demand strict formatting.
3. The hallucination rate generally remains low, especially when increasing the number of epochs, which we speculate is due to the relative simplicity of our dataset. Since most of our data provides a clear context and our extracted description does not involve complex language structures so that the model is more likely to produce responses within the whole corpus.
4. While recall gradually increases with the number of epochs, the overall recall values remain modest. We hypothesize this is due to an abundance of selectable APIs for some queries. Despite limiting identical function APIs, in cases like image classification, common object labels appearing across multiple datasets can accumulate suitable APIs, potentially exceeding ten options. Therefore, even with ten inference times, it’s impossible to cover all possibilities.
Overall, the results affirm that our system Tool-LMM effectively brings out the tool-use potential in LLMs and enhances the LLMs to deal with multi-modal instruction inputs.
Model Video Audio Image Text
[…]
Vicuna-7B (3epochs) 76.19 4.76 88.41 3.05 79.87 1.18 73.62 3.58
Llama-7B (3epochs) 85.71 0.00 92.68 0.00 75.66 1.18 76.38 6.20
Llama2-7B (3epochs) 80.95 0.00 89.63 0.61 76.97 1.05 72.18 4.61
Llama2-Chat-7B (3epochs) 76.19 0.00 87.20 0.00 79.87 0.66 73.00 7.78
Vicuna-7B (5epochs) 71.43 0.00 93.29 0.00 88.95 0.53 79.13 2.41
Llama-7B (5epochs) 76.19 0.00 89.02 0.00 87.24 0.26 82.64 0.83
Llama2-7B (5epochs) 80.95 0.00 93.29 0.00 83.82 0.39 77.62 3.03
Llama2-Chat-7B (5epochs) 80.95 0.00 92.07 0.00 86.45 1.84 81.20 1.24
Vicuna-13B (3epochs) 85.71 0.00 89.63 1.83 76.71 1.84 72.59 4.13
Llama-13B (3epochs) 80.95 0.00 92.07 0.00 72.50 1.45 73.21 5.23
Llama2-13B (3epochs) 76.19 0.00 89.63 0.00 78.16 1.32 75.83 4.89
Llama2-Chat-13B (3epochs) 85.71 0.00 90.24 0.00 44.61 1.32 72.04 4.89
Vicuna-13B (5epochs) 80.95 0.00 89.63 0.00 86.18 1.97 78.03 2.27
Llama-13B (5epochs) 76.19 0.00 90.85 0.00 79.87 0.92 80.30 2.20
Llama2-13B (5epochs) 71.43 0.00 92.07 0.00 82.11 0.13 79.89 2.27
Llama2-Chat-13B (5epochs) 76.19 0.00 85.37 0.00 44.21 0.53 77.69 1.93
Vicuna-13B (10epochs) 76.19 0.00 92.07 0.00 90.66 0.13 86.64 0.21
Llama-13B (10epochs) 71.43 0.00 92.07 0.00 92.76 0.00 85.06 0.28
Llama2-13B (10epochs) 85.71 0.00 89.02 0.00 71.84 0.13 86.16 0.21
Llama2-Chat-13B (10epochs) 80.95 0.00 89.63 0.00 90.53 0.00 86.23 0.14
Table 8: The performance on testing subsets divided by input modality.
6.4 Results Analysis
To verify the effectiveness of our system when encountering different complex cases, we conduct experiments on all types of testing subsets defined by three previously established criteria in Sec. 5, assessing the impact of different characteristics of data on our model’s performance and accuracy. All the 7B models are trained in 5 epochs and 13B models are trained in 10 epochs by default for a fair comparison.
Ambiguity types. In Table 6, we provide the results for five types of ambiguity, and the comparison between non-ambiguity and ambiguity cases. We find that the ambiguity types I(Domains) and II(Categories) perform worst in ambiguity handling. In contrast, other categories excel, especially type V(Language) achieves 100% accuracy when trained 5 epochs for 7B or 10 epochs for 13B models. Intuitively, the exceptional performance of type III(Quality) and type IV(Conditions) may stem from significant visual differences and clear classification boundaries. In contrast, for type I and type II, particularly the latter, the challenges arise due to varying granularity in datasets. The model must not only understand the multimodal input content but also discern the most appropriate classification granularity for the task at hand, complicating the choice of API. For the overall performance, we find that performance without ambiguity surprisingly falls short compared to that with ambiguity. It indirectly highlights that text descriptions cannot substitute for multimodal information, even when the language is exceptionally detailed. This underscores the necessity for LLMs to process the capability to perceive visual- and auditory- instructions, demonstrating the significance of our work.
Model one-to-one one-to-many
[…]
Vicuna-7B (3epochs) 51.75 6.71 86.62 1.23
Llama-7B (3epochs) 56.12 6.85 85.86 3.04
Llama2-7B (3epochs) 47.67 6.56 85.91 1.81
Llama2-Chat-7B (3epochs) 57.29 6.56 83.75 4.27
Vicuna-7B (5epochs) 66.03 3.35 90.01 0.94
Llama-7B (5epochs) 69.24 1.31 90.59 0.29
Llama2-7B (5epochs) 63.12 3.64 87.73 1.29
Llama2-Chat-7B (5epochs) 70.99 2.33 88.66 0.94
Vicuna-13B (3epochs) 47.96 7.14 86.09 1.64
Llama-13B (3epochs) 47.23 5.69 85.21 2.81
Llama2-13B (3epochs) 51.90 8.16 87.78 1.46
Llama2-Chat-13B (3epochs) 41.84 7.00 73.87 1.93
Vicuna-13B (5epochs) 61.95 4.52 89.25 0.99
Llama-13B (5epochs) 61.81 3.35 88.49 0.94
Llama2-13B (5epochs) 62.97 3.06 88.72 0.76
Llama2-Chat-13B (5epochs) 50.58 3.35 74.40 0.53
Vicuna-13B (10epochs) 77.26 0.58 92.58 0.00
Llama-13B (10epochs) 76.68 0.58 92.34 0.00
Llama2-13B (10epochs) 69.24 0.44 86.85 0.06
Llama2-Chat-13B (10epochs) 76.38 0.29 92.34 0.00
Table 9: The performance on testing subsets divided by API option number.
Multiple options. We divide the testing set into two to distinguish whether prompts can be linked to multiple API options(one-to-one or one-to-many), and the corresponding performances are shown in Table 9. Surprisingly, we discover that one-to-one performance does not surpass, and is significantly behind, the scenario with one-to-many cases, lagging by an average of 21.90% across four 7B models at 5 epochs and 16.14% across four 13B models at 10 epochs . Besides, in terms of the rate of hallucinations, the one-to-one demonstrates a relatively higher incidence. Our analysis suggests that one-to-one pairs may lead to overfitting, where the model becomes overly tuned to specific pairs, hindering its ability to generalize to slightly different queries or contexts. The higher rate of hallucinations also verifies the above speculation.
Individual modality. We further investigate the model’s perception and understanding of instructions with the specific sole modality and report the results in Table 8. Notably, the audio modality stands out, achieving an impressive performance among all models, reaching as high as 93.29% on both the Llama2-Chat-7B and Vicuna-7B model. On the other hand, we notice that pure text-based performance is not particularly satisfactory, falling short in both accuracy and hallucination rate. As for the video modal, it has a greater performance variance than other modals. Interestingly, We find that as the number of epochs increased, the performance of Vicuna-7B drops significantly. A similar situation also occurs on the several other models. We speculate that this may be related to the complexity and scarcity of video data. Since our video task mostly involves video understanding, but with the epochs increasing, the model might start to emphasize the detailed information, which can impact the performance when extending to new unseen data.
Figure 5: Qualitative example of Tool_LMM on text ambiguity case. Tool_LMM firstly receives the text instruction and perceives information from the input image to determine the corresponding task, select the appropriate API that meets the requirements from the API corpus, and then call the API to execute, and finally get the quality-improved image. Zoom in for the best view.
6.5 Visualization
We visualize qualitative result of Tool_LMM, to verify its effectiveness. As shown in Figure 5, our system is asked to improve the quality of the input image without expaining the reason for the poor image quality, and Tool_LMM is capable of perceiving the multi-modal information and choosing the correct API by combining the instruction and visual information, which proves that our Tool_LMM can effectively accomplish the users’ needs.
7 Conclusion
In this work, we introduce how to empower the tool-augmented LLMs with perceiving visual- or audio-grounded instructions’ information. We collect a benchmark, ToolMMBench, from the HuggingFace Platform, which covers 932 high-quality machine learning related APIs. The finetuning of this dataset on various popular LLMs has yielded promising results, demonstrating that Tool-LMM can accurately recommend appropriate tools for multi-modal instructions.
[…]
Supplementary Material
In this supplementary material, we list i) dataset details, including selection criteria, prompt templates, data distribution; ii) our proposed categorization system.
8 Dataset Details
In this section, we elaborate on the details of dataset construction and the distribution of training and testing data. Specifically, Figure 8.1 visualizes the entire dataset construction process, while Sec. 8.2 - Sec. 8.4 detail each step, including API selection criteria, API functional boundaries criteria, and prompt construction. Sec. 8.5 visualizes some data distributions of the training set and testing set. We emphasize that the collection and processing of this dataset involved over 2000 person-hours, ensuring its reliability for advanced computer vision research and applications by meticulous attention to detail.
Figure 6: The process of dataset construction. It mainly consists of three stages. The first step is collection and preprocessing, including crawling the original API from HuggingFace, filtering and re-categorizing the API according to tasks. The second step is the generation of instruction-answer pairs, including using GPT-4 to generate instructions based on API functions, determining whether there is text ambiguity in the user prompts, and searching for content-matching input for multi-modal. The third step is to construct a potential one-to-many instruction-answer pair based on the relationship between API functions.
Figure 7: Five types of ambiguity cases needed to be checked during the instruction-answer pair generation step.
8.1 Visualization of Dataset Construction Process
We show the whole process of collecting our dataset in Figure 6. The difficulty in collecting the benchmark is reflected in three aspects. (1) Task Classification: We restructure a hierarchical task classification system. This advancement rectifies the overly coarse-grained categorizations observed in the HuggingFace platform, ensuring that each API is endowed with a specific and unambiguous task label. (2) Annotation of Multimodal Input Instructions: Should an instruction reveal cues from modalities other than text, a precise coherence between the divulged information in the instruction and the content of the non-textual modality is paramount. In cases where no such cues are disclosed, we deem it essential to investigate the presence of five potential cases that are caused by textual ambiguity. And we show the visualization of five ambiguity types in Figure 7. Such scrutiny lays the groundwork for subsequent experiments aimed at validating whether information from other modalities aids large language models in making appropriate tool selections; (3) Instruction Matching: To identify all APIs meeting an instruction’s functional requirements, we must categorize APIs based on identical and similar functionalities respectively during collection. When pairing instructions for groups of APIs with analogous functions, it becomes crucial to discern whether a provided instruction strictly satisfies each API’s functional description. This determination will influence the number of APIs that can aptly respond as the correct answer to a given instruction.
8.2 Filtering Rules
After crawling APIs from HuggingFace, due to varying quality among these models, we implement stringent selection criteria for APIs. The specific filtering rules for this selection process are outlined as follows:
Low-Quality Model Card A significant number of APIs either lack model cards or contain overly simplistic ones, providing little helpful information. Additionally, some APIs have ceased maintenance or have been integrated into other APIs. Even multiple APIs share a single model card. These APIs pose challenges in organizing API documentation and require manual consultation of additional sources like papers and GitHub. While enriching the number of APIs, this approach introduces inaccuracies. Finally, we opted to remove such APIs from our dataset uniformly.
NSFW Content Risk For ethical and safety considerations, we remove APIs, particularly some of the text-to-image type, which pose risks of generating NSFW (Not Safe For Work) content. Such content is deemed inappropriate for all age groups and could compromise the dataset’s applicability in diverse settings.
License Restrictions We exclude models with restrictive licensing terms. An example is the Llama2  [34] model, which requires a formal application and approval process to download and use its model weights.
Identical APIs Numerous models on HuggingFace share identical structures and datasets, merely differing in implementation by various contributors. To avoid redundancy, we selectively retain only the most downloaded version of each such model, ensuring the dataset is more streamlined. For APIs with different model architectures but trained on the same dataset and having the same functions, as we mentioned in the main text, we select the top five APIs based on download statistics.
8.3 API Function Boundaries
Multi-task APIs Our dataset includes multi-task APIs like various large language models (Bloom  [36], Baichuan2  [40], Falcon  [22]), capable of numerous NLP tasks, some of which even fall outside HuggingFace’s classification system. For simplicity, we focus specifically on their text generation capabilities, aligning with HuggingFace’s functional categorization of such APIs. Additionally, we modify the descriptions of these APIs in our dataset to ensure a match between the functionalities reflected by the provided queries and those described in the API descriptions.
API Function Differentiation After identifying the specific subdomain of each API, we detail our approach for discerning whether APIs within the same subdomain possess identical or similar functions. For plain text input APIs, we can assess APIs based on the language and context of the datasets they utilize. For example, in sentiment analysis tasks, we discern various contexts by relying on the distinct domains of data sources, such as financial, legal, social media, and reviews, enabling us to classify nuanced functional differences among similar-function APIs. Similarly, for multimodal inputs, especially images, we refine API function categorization by analyzing dataset contexts. For instance, in Visual Question Answering tasks, datasets feature a range of images such as Diagrams, Charts, Documents, Infographics, and other general images, enabling nuanced API segmentation. The above examples provide merely a subset of our API classification rules. Due to the complex and case-by-case nature of the analysis, the complete set of rules is not listed here.
One-to-many Situation Identification Originating from each API’s functional description, our approach, driven by GPT-4, generates corresponding instructions that adhere strictly to the API’s functional boundaries. Hence, consideration is primarily given to other APIs with the same function in one-to-many scenarios.
8.4 Prompt Construction
Below, we list the detailed prompts used for generating instructions. The prompts fed to GPT-4 are composed of (1)the API call’s relevant information, including its name, a concise functional description, and training datasets if mentioned in the model card, (2)some sample instructions, and (3)additional requirements. The sample instructions are leveraged to foster format standardization and enhance the quality of GPT-4’s outputs. So we invite some expert annotators to craft some high-quality instructions in advance and then randomly choose two each time as exemplary inputs to GPT-4. We design two sets of prompt templates in our dataset, mindful of prompt ambiguity cases. For APIs with multimodal inputs susceptible to prompt ambiguity, we strive to balance the final selection of ten instructions, maintaining a 1:1 ratio between unambiguous and ambiguous instructions; the corresponding prompt template is shown in Figure 8 and Figure 9 respectively. We emphasize that ambiguity pairs in queries are those pairs with identical textual inputs but varied multimodal inputs.
### User: You are an NLP task expert. Given an API, you need to generate 20 different user queries that utilize the API function, adhering to the following input and output format for each query: Input: API Name: This is the name of the API Function. Description: This is a detailed description of the model. Prohibit Words: These are words that are prohibited from appearing in the output. Output: Query1: This is an instruction that can utilize the API function. Query2: This is an instruction that can utilize the API function. Query3: This is an instruction that can utilize the API function. Query4: This is an instruction that can utilize the API function. Query5: This is an instruction that can utilize the API function. Query20: This is an instruction that can utilize the API function. Below are some examples: Example 1: Input: API Name: microsoft/codereviewer Description: CodeReviewer is a model pre-trained with code change and code review data to support code review tasks. Prohibit Words: ”API, tools, model” Output: Query1: I’ve made some changes to my Python code. Can you review this snippet for me? ”def addNums(a, b): return a + b” Query2: I’m not very confident in my Java coding skills. Could you check this piece for potential issues? ”public int divide(int x, int y) { return x / y; }” Query3: I’m trying to refactor this C# method. Does it look okay? ”public void PrintName(string name) { Console.WriteLine(name); }” Query4: I wrote this SQL query, and I’m unsure if it’s optimized. Can you review it? ”SELECT * FROM users WHERE age $\geq$ 25;” Query5: Here’s a piece of JavaScript function I came up with. Any suggestions for improvement? ”function greet(name) { return ’Hello ’ + name; }” Query6: I’ve been learning Ruby recently. Would appreciate your thoughts on this code segment: ”def multiply(x, y) x * y end” Query7: I’m not sure if this CSS is correct. Can you review it? ”h1 { color: blue; font-size: 30px; }” Query8: I just started with Swift. Here’s a function I wrote. Could you review it for best practices? ”func displayMessage(message: String) { print(message) }” Query9: My colleague wrote this PHP function, but it seems off. Can you review it for me? ”function subtract(x,x,y) { return $x$-y; }” Query20: I’m learning Kotlin and wrote this simple method. Mind taking a look? ”fun sum(a: Int, b: Int): Int = a + b” Example 2: Input: API Name: Salesforce/blip-vqa-base To be continued
Continued
Description: This model is a base-sized ViT-based BLIP trained on the visual question answering task VQA2.0 dataset. The model’s input is an image and a text question, and the model’s output is a text answer.
Prohibit Words: ”API, tools, model, VQA2.0” Output: Query1: I recently visited a new city and took some photographs. Based on this particular image, can you tell me what is the weather conditions like? Query2: I was at a photo exhibition and took this picture. Could you provide information about was this taken in a Latin American country judging from the dress? Query3: I was at the park and saw a dog playing. By analyzing this photo, can you tell if the frisbee it is using is hard or soft? Query4: My cousin went to a new food joint and ordered their specialty hot dog. By looking at this picture, can you determine if the hot dog is larger than a normal one? Query5: I attended a family gathering and took a group photo. By analyzing the photograph, how many kids can you see in the picture? Query6: Looking at this captivating portrait, I noticed an object in the individual’s hand. It seems to be a sweet treat. What is this person holding? Query7: I was at the beach and took a photo of a man skiing. By examining the image, can you tell if the wave is chasing him? Query8: I visited the local train station and captured an image of a passing train. Can you identify what color the train is from the picture? Query9: This cityscape presents an old-fashioned vehicle that stands out amidst the modern structures. I’d like to know more about this vehicle. Could you tell me its model and color?” Query20: Given the serene backdrop of sailboats and a calm sea, can you tell me how many boats are there? Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
Figure 8: Example of the unambiguous prompt template: This template applies to pure text inputs and a part of multimodal input scenarios where textual inputs reveal multimodal information. In practical usage, the template’s masked sections should be populated with the respective API’s name, description, and words prohibited in the instructions.
### User: You are an NLP task expert. Given an API, you need to generate 20 different user queries that utilize the API function, adhering to the following input and output format for each query: Input: API Name: This is the name of the API Function. Description: This is a detailed description of the model. Prohibit Words: These are words that are prohibited from appearing in the output. Output: Query1: This is an instruction that can utilize the API function. Query2: This is an instruction that can utilize the API function. Query3: This is an instruction that can utilize the API function. Query4: This is an instruction that can utilize the API function. Query5: This is an instruction that can utilize the API function. Query20: This is an instruction that can utilize the API function. Below are some examples: Example 1: Input: API Name: timm/resnet101.a1h_in1k Description: A 101 layers ResNet-B image classification model trained on ImageNet-1k. Prohibit Words: ”API, tools, model, ImageNet” Output: Query1: I’d value your input on this image’s classification. Query2: There’s an image in my possession, and I’m seeking a label for it. Query3: I need help in determining the label for this image. Can you help? Query4: Can you provide a label perspective for this image? Query5: How would you manage the categorization of this particular visual? Query6: There’s an intriguing image that I’ve come across, and I’m curious about its category. Query7: How would you classify this picture? Query8: I’d be grateful for your view on this image’s category. Query9: I’m searching for a label for this visual. Can you assist? Query20: What category seems plausible for this visual, in your view? Example 2: Input: API Name: lllyasviel/control_v11f1p_sd15_depth Description: This model is intended for control diffusion models by adding extra conditions. Trained with depth estimation, the condition image is an image with depth information, usually represented as a grayscale image, and the output is a new image. Prohibit Words: ”API, tools, model, depth, controlnet” Output: Query1: Use my image to depict a family picnic in a sunlit park with children playing. Query2: Can you reimagine my image as a soccer match in progress, with players and fans cheering? Query3: Create an adventure scene with pirates and treasure islands using my image. Query4: Can you make my image into a heartwarming scene of a mother duck leading her ducklings through a pond? To be continued
Continued
Query5: Craft a depiction of a magician performing tricks in front of an amazed audience from my image.
Query6: Could you reshape this photograph into a bustling train station with intricate ironwork and arches? Query7: Transform this snapshot into a charming cottage nestled among rolling hills and wildflowers. Query8: I’m curious about how this picture would be reimagined as a bustling harbor with ships, cranes, and cargo containers. Query9: Use my image to depict a family picnic in a sunlit park with children playing. Query20: I’d love to see this image as a gardener carefully pruning roses in a blooming garden. Note that: 1. When crafting queries, avoid including the API’s name; 2. Ensure that the queries are varied and diverse; 3. When processing an input, any words listed in the Prohibited Words must be strictly excluded from the response; 4. The queries should not convey or imply multimodal information. Now, let’s start. Input: API Name: [API Name] Description: [API Description] Prohibit Words: [Some words]
Figure 9: Example of the ambiguous prompt template: This template applies to other multimodal input scenarios where textual inputs do not reveal multimodal information. In practical usage, the template’s masked sections should be populated with the respective API’s name, description, and words prohibited in the instructions.
8.5 Visualization of Data Distributions
In our work, we collect 932 APIs and one ”Unknown” category, featuring 651 APIs with pure text inputs, 191 APIs incorporating image inputs, 80 incorporating audio inputs, and 10 incorporating video inputs. Given the long-tail distribution of API quantities for various tasks on HuggingFace, coupled with our selection criteria, the APIs we ultimately choose exhibit a similar long-tail distribution across both modality and task level. In future work, we aim to enrich our collection with more APIs, particularly those with video inputs, and optimize the distribution across different tasks for more uniformity.
9 Categorization System
We illustrate our proposed categorization system in Table 10. The system is two-level hierarchical. The coarse level task setting mainly refers to the classification of HuggingFace, while the fine-grained subtask is based on the different task descriptions.
Task Subtask Api Names
Audio Classification Event Recognition MIT/ast-finetuned-audioset-10-10-0.4593
Command Recognition MIT/ast-finetuned-speech-commands-v2
Spoken Language Identification TalTechNLP/voxlingua107-epaca-tdnn
Speaker Verification anton-l/wav2vec2-base-superb-sv
Emotion Recognition audeering/wav2vec2-large-robust-12-ft-emotion-msp-dim
Gender Recognition m3hrdadfi/hubert-base-persian-speech-gender-recognition
Keyword Spotting superb/hubert-base-superb-ks
Audio-to-Audio Single-Channel Speech Enhancement JorisCos/ConvTasNet_Libri1Mix_enhsingle_16k
Speech-to-Speech Translation facebook/xm_transformer_unity_en-hk
Voice Conversion microsoft/speecht5_vc
Seperate Clean mpariente/DPRNNTasNet-ks2_WHAM_sepclean
Audio Source Separation speechbrain/sepformer-libri3mix
Pop-to-Piano sweetcocoa/pop2piano
Feature Extraction Audio Feature Extraction LeBenchmark/wav2vec2-FR-7K-large
Audio-to-Text Automatic Speech Recognition Harveenchadha/vakyansh-wav2vec2-hindi-him-4200
Voice Activity Detection Speaker Segmentation philschmid/pyannote-segmentation
Speaker Diarization philschmid/pyannote-speaker-diarization-endpoint
Depth Estimation Depth Estimation Intel/dpt-hybrid-midas
Image Classification General Image Classification facebook/convnextv2-tiny-1k-224
Specific Type Image Classification Neto71/sea_mammals
Style Classifier playrobin/furniture-styles
Image Segmentation Text Guided Image Segmentation CIDAS/clipseg-rd64-refined
Semantic Segmentation apple/deeplabv3-mobilevit-small
Panoptic Segmentation facebook/detr-resnet-50-panoptic
Instance Segmentation facebook/mask2former-swin-small-coco-instance
Zero-shot Segmentation facebook/sam-vit-huge
Image-to-Image Image Super-Resolution caidas/swin2SR-classical-sr-x2-64
Image Variations lambdalabs/sd-image-variations-diffusers
2D-to-3D openai/shap-e-img2img
Image Deblurring google/maxim-s3-deblurring-gopro
Image Dehazing google/maxim-s2-dehazing-sots-outdoor
Image Deraining google/maxim-s2-deraining-rain13k
Feaure Extraction Image Feature Extraction BridgeTower/bridgetower-base
Object Detection Object Detection SenseTime/deformable-detr
Table Detection TahaDouaji/detr-doc-table-detection
Text-conditioned Object Detection google/owlvit-base-patch32
Visual Question Answering Visual Question Answering Salesforce/blip-vqa-base
Chart Question Answering google/matcha-chartqa
Diagram Question Answering google/pix2struct-ai2d-base
Document Question Answering google/pix2struct-docvqa-base
Infographics Question Answering google/pix2struct-infographics-vqa-large
Task Subtask Api Names
Image-to-Text Conditional Image Caption Salesforce/blip2-flan-t5-xl
Optical Character Recognition alibaba-damo/mgp-str-base
Image Caption bipin/image-caption-generator
Document Parsing naver-clova-ix/donut-base-finetuned-cord-v2
Tag Generation SmilingWolf/wd-v1-4-convnextv2-tagger-v2
Chart-to-Table google/deplot
Zero-Shot Image Classification Zero-Shot Image Classification laion/CLIP-ViT-B-16-laion2B-s34B-b88K
Multimodal-to-Image Face-Detection-Conditioned Prompt-guided Image Generation CrucibleAI/ControlNetMediaPipeFace
Prompt-guided Cartoonization instruction-tuning-sd/cartoonizer
Brightness-Conditioned Prompt-guided Image Generation ioclab/control_v1p_sd15_brightness
Prompt-guided Pixel-to-Pixel Image Editing lllyasviel/control_v11e_sd15_ip2p
Shuffle-Conditioned Prompt-guided Image Generation lllyasviel/control_v11e_sd15_shuffle
Tiled-Conditioned Prompt-guided Image Generation lllyasviel/control_v11f1e_sd15_tile
Depth-Conditioned Prompt-guided Image Generation lllyasviel/control_v11f1p_sd15_depth
Canny-Edges-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_canny
Line-Art-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_lineart
MLSD-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_mlsd
Normal-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_normalbae
Openpose-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_openpose
Scribble-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_scribble
Segment-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_seg
HED-Boundary-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15_softedge
Anime-Line-Art-Conditioned Prompt-guided Image Generation lllyasviel/control_v11p_sd15s2_lineart_anime
Prompt-guided Image Variations stabilityai/stable-diffusion-2-1-unclip
Text Classification Sentiment Analysis DTAI-KULeuven/robbert-v2-dutch-sentiment
Emotion Analysis MilaNLProc/feel-it-italian-emotion
Offensive Language Detection Hate-speech-CNERG/bert-base-uncased-hatexplain
Topic Classification MaartenGr/BERTopic_Wikipedia
Toxicity Analysis OpenAssistant/reward-model-deberta-v3-large-v2
Sensitive Text Detection apanc/russian-inappropriate-messages
Irony Detection cardiffnlp/twitter-roberta-base-irony
Bias Detection cffl/bert-base-styleclassification-subjective-neutral
Feature Extraction Text Feature Extraction facebook/bart-large
Fill-Mask Fill-Mask CLTL/MedRoBERTa.nl
Text Generation Text Generation EleutherAI/gpt-j-6b
Prompt Generation FredZhang7/anime-anything-promptgen-v2
Code Generation NumbersStation/nsql-350M
Specific Text Genre Generation uer/gpt2-chinese-poem
Task Subtask Api Names
Text Generation Text Generation EleutherAI/gpt-j-6b
Prompt Generation FredZhang7/anime-anything-promptgen-v2
Code Generation NumbersStation/nsql-350M
Specific Text Genre Generation uer/gpt2-chinese-poem
Question Answering Extractive Question Answering CATIE-AQ/QAmembert
Open Domain Question Answering facebook/dpr-ctx_encoder-single-nq-base
Sentence Similarity Sentence Similarity intfloat/e5-small-v2
Summarization Summarization IDEA-CCNL/Randeng-Pegasus-523M-Summary-Chinese
Title Generation JulesBelveze/t5-small-headline-generator
Keyword Generation Voicelab/vlt5-base-keywords
Text-to-Image Text to General Style Image runwayml/stable-diffusion-v1-5
Text to RGBD Image Intel/ldm3d
Text to Specific Style Image nitrosocke/spider-verse-diffusion
Text to 3D Image openai/shap-e
Text to Spectrogram Image riffusion/riffusion-model-v1
Text-to-Speech Text-to-Speech Voicemod/fastspeech2-en-male1
Text-to-Video Text-to-Video damo-vilab/text-to-video-ms-1.7b
Text-to-Text General Text Generation 1-800-BAD-CODE/xlm-roberta_punctuation_fullstop_truecase
Sentence Correction KES/T5-KES
Generative Question Answering MaRiOrOsSi/t5-base-finetuned-question-answering
Question Generation allenai/t5-small-squad2-question-generation
Paraphraser cointegrated/rut5-base-paraphraser
Recipe Generation flax-community/t5-recipe-generation
Question Answering Generation google/t5-small-ssm-nq
Text Revision grammarly/coedit-large
Relation Extraction ibm/knowgl-large
Code Review microsoft/codereviewer
Commonsense Reasoning mrm8488/t5-base-finetuned-common_gen
Span Sentiment Extraction mrm8488/t5-base-finetuned-span-sentiment-extraction
Distractor potsawee/t5-large-generation-race-Distractor
Detoxification s-nlp/bart-base-detox
Symbolic Music Generation sander-wood/text-to-music
Translation Translation Babelscape/mrebel-large
Zero-Shot Classification Zero-Shot Classification MoritzLaurer/DeBERTa-v3-base-mnli-fever-anli
Video Classification Human action recognition video classification MCG-NJU/videomae-base-finetuned-kinetics
Video-to-Text Video Question Answering kpyu/video-blip-flan-t5-xl-ego4d
Video caption Neleac/timesformer-gpt2-video-captioning
Feature Extraction Video Feature Extraction deepmind/multimodal-perceiver
Table 10: Our proposed categorization system. The first and second columns are the two levels of the categories for each task. The third column provides an example of the candidate models.